OpenStudy (anonymous):

Am I right? Find the derivative of f(x) = 6 divided by x at x = -2. The answer would be -(3/2)

OpenStudy (solomonzelman):

\[f(x)=\frac{6}{x}\] like this ?

OpenStudy (anonymous):

Mhm. I think it's either that or just 6.

OpenStudy (anonymous):

So.. Am I right or wrong?

OpenStudy (anonymous):

f(x)= 6/x my answers are either -(3/2) or 6.

OpenStudy (anonymous):

@IAmSinged

OpenStudy (anonymous):

Put the equation as 6x^-1 since it's easier to use the product rule on if you're new to calculus. The derivative of that is -6x^(-2). Plugging in -2 gets me 3/2.

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

*-3/2

OpenStudy (anonymous):

oh! ohkay, so I was right. thank you!

OpenStudy (anonymous):

what if I have of f(x) = -6/x at x = 12.

OpenStudy (anonymous):

it would be 1/2, right?

OpenStudy (anonymous):

@IAmSinged

OpenStudy (anonymous):

If you want f'(x) at 12: f(x) = -6x^-1 f'(x) = 6x^-2 = 6/x^2 f'(12) = 6/144 = ~0.04167

OpenStudy (anonymous):

1/24 then?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

Thank you very much then. c: